Ivanhoe Corporation purchased a new machine for its assembly process on August 1, 2017. The cost of this machine was $146,196. The company estimated that the machine would have a salvage value of $15,996 at the end of its service life. Its life is estimated at 5 years, and its working hours are estimated at 21,200 hours. Year-end is December 31 Compute the depredation expense under the following methods. Each of the following should be considered unrelated. (Round depreciation rate per hour to 2 decimai places, e.g. 5.35 for computational purposes. Round your answers to o decimal places, e.g. 45,892.) (a) Straight-line depreciation for 2017 (b) Activity method for 2017, assuning that machine usage was 810 hours c) Sum-of-the-years-digits for 2018 4974.62 )Double-decining balance for 2018 48732
Answer #1

a)           

Straight line depreciation = (Cost of asset – Salvage value)/ useful life

Straight line depreciation expense for 2017 = [($146196 - $15996)/5]*5/12

= ($130200/5)*5/12 = $26040*5/12 = $10850.

b)

Activity method = (Cost of asset – Salvage value)*Current year usage / Total estimated working hours

= ($146196 - $15996)*810/21200 = $130200*810/21200 = $4974.62.

c)

Sum of year digit methods = n (n+1)/2 = 5(5+1)/2 = 5*3 = 15

Where n = numbers of useful life

Remaining useful life in year 2018 = Useful life – year of asset already used = 5-1 = 4 years

Depreciation expense in year 2018 = First seventh months i.e. up to July 31 2018 =

(Cost of asset – Salvage value)*Remaining useful life / sum of year digit

[($146196 - $15996)*5/15]*7/12 = ($130200*5/15)*7/12 = $43400*7/12 = $25316.67

Next 5 month (From August 01 – December 31) =

[($146196 - $15996)*4/15]*5/12 = = ($130200*4/15)*5/12 = $34720*5/12 = $14466.67

Total Depreciation expense by using sum of year digit methods in year 2018 =

$25316.67 + $14466.67 = $39783.34. i.e. $39783

d)

Double declining balance methods = Cost of asset * Double declining rate

Double declining rate = (100/useful life)*2 = (100/5)*2 = 40%

Depreciation expense in 2017 = ($146196*40%)*5/12 = $58478.4*5/12 = $24366

Depreciation expense in 2018 = ($146196 -$24366)*40% =$121830*40% = $48732